he current in a tungsten filament lamp is 0.25 A. The filament has a radius o


f 5.4 × 10


−6


m


2


.


Tungsten has 3.4 × 10


28


electrons per cubic metre. Show that the mean drift velocity of the


electrons in this very thin filament is about 0.5 m s


−1



1
Expert's answer
2021-11-10T10:08:52-0500

By the definition of the drift velocity, we have:


"v_d=\dfrac{I}{nAq}=\dfrac{I}{n\pi r^2q},""v_d=\dfrac{0.25\ A}{3.4\times10^{28}\ \dfrac{el}{m^3}\times\pi\times(5.4\times10^{-6}\ m)^2\times1.6\times10^{-19}\ C},""v_d=0.5\ \dfrac{m}{s}."